OpenStudy (anonymous):

Find f'(a) f(x)=square root of 1-2x using the equation:limit as x approaches 0 f(a+h)-f(a)/h

zepdrix (zepdrix):

\[\Large f(\color{#DD4747 }{x})\quad=\quad \sqrt{1-2(\color{#DD4747 }{x})}\] \[\Large f(\color{#DD4747 }{a})\quad=\quad \sqrt{1-2(\color{#DD4747 }{a})}\]\[\Large f(\color{#DD4747 }{a+h})\quad=\quad \sqrt{1-2(\color{#DD4747 }{a+h})}\] Understand how function notation works? We can now use these two pieces to plug them into our difference quotient.\[\Large \lim_{h\to0}\frac{f(a+h)-f(a)}{h}\]

OpenStudy (anonymous):

I see some parts of it...did you send me a link? It looks like it but Im not sure?

zepdrix (zepdrix):

A link? No. Are the math equations not showing up? Hmm you are using Internet Explorer I assume? :(

zepdrix (zepdrix):

The math plugin doesn't work with IE :c

OpenStudy (anonymous):

Oh ok will Mozilla fox work?

OpenStudy (isaiah.feynman):

Wait, what is 'a' itself?

OpenStudy (anonymous):

The derivative srry : (

OpenStudy (anonymous):

Ok I see the picture now. I got to that point finally. What is the next thing I need to do?

zepdrix (zepdrix):

So we plug in our pieces: \[\Large \lim_{h\to0}\frac{\sqrt{1-2(\color{#DD4747 }{a+h})}-\sqrt{1-2(\color{#DD4747 }{a})}}{h}\]

zepdrix (zepdrix):

Then we want to multiply the top and bottom by the `conjugate of the top`.

zepdrix (zepdrix):

So we want to multiply the top and bottom by:\[\Large \left(\frac{\sqrt{1-2(\color{#DD4747 }{a+h})}+\sqrt{1-2(\color{#DD4747 }{a})}}{\sqrt{1-2(\color{#DD4747 }{a+h})}+\sqrt{1-2(\color{#DD4747 }{a})}}\right)\]

OpenStudy (anonymous):

Ok I see that part there. So next we simplify or simplify it further?

zepdrix (zepdrix):

Don't touch the denominator! But yes we can simplify the numerator `quite a bit`.

OpenStudy (anonymous):

Do you get 2h for the numerator?

zepdrix (zepdrix):

Mmm yah that sounds right so far!

OpenStudy (anonymous):

Well I probably got it the wrong way so please let me know...So I got -2h/h times square root 1-2a-2h + square root of 1-2a then I added up like terms and got -2h/h times square root of 2-4a-2h...then I just took out like terms or simplified it to -1/square root 1-2a...now I looked in the back of the book but I wanted to make sure I got it the right way..

zepdrix (zepdrix):

Mmm I'm not quite sure what you're saying :( The 2h sounded right, leaving you here,\[\Large \lim_{h\to0}\frac{2h}{h\left(\sqrt{1-2(\color{#DD4747 }{a+h})}+\sqrt{1-2(\color{#DD4747 }{a})}\right)}\]The h's cancel,\[\Large \lim_{h\to0}\frac{2}{\left(\sqrt{1-2(\color{#DD4747 }{a+h})}+\sqrt{1-2(\color{#DD4747 }{a})}\right)}\] And now it's in a nice form where we can simply plug h=0 directly in.

zepdrix (zepdrix):

Oh I missed a minus sign, woops\[\Large \lim_{h\to0}\frac{-2}{\left(\sqrt{1-2(\color{#DD4747 }{a+h})}+\sqrt{1-2(\color{#DD4747 }{a})}\right)}\]

zepdrix (zepdrix):

It sounds like you got to the answer, or no? I can't tell based on the way you described it :(

OpenStudy (anonymous):

Yes but you explained it in a way that I'll copy the way you did it instead.haha

zepdrix (zepdrix):

After you get your -2h on top, you cancel out an h from the top and bottom. Then we plug in h=0,\[\Large \frac{-2}{\left(\sqrt{1-2(\color{#DD4747 }{a})}+\sqrt{1-2(\color{#DD4747 }{a})}\right)}\]And we can add the sqrt's,\[\Large \frac{-2}{2\sqrt{1-2a}}\]And then simplify a little further.

OpenStudy (anonymous):

Ok I see it now thank u